# Global AI Governance Overview: Understanding Regulatory Requirements Across Global Jurisdictions

### v1

@@ 3.8.3 South Korea
South Korea became the first jurisdiction outside the EU to enact a comprehensive AI law, the Basic Act on the Development of AI and the Establishment of Trust (“AI Basic Act”), officially passed on December 26, 2024, and slated to take effect in January 2026, following a one-year preparation period [korea2024AIBasicAct]. The Act focuses on “high-impact AI” systems deployed in critical areas such as healthcare, education, finance, employment, and essential public services. These systems must be designed to allow continuous, meaningful human oversight and intervention, and providers are required to notify users if they are interacting with AI-generated content or decisions, conduct risk assessments, maintain proper documentation, and implement effective control measures [korea2024AIBasicAct]. Additionally, foreign AI providers must appoint a local representative to handle regulatory communications [korea2024AIBasicAct].
This new law operates alongside existing statutes: the Personal Information Protection Act (PIPA), which mandates user consent and restricts the scope of personal data collection [koreaPIPA]; the Network Act (Act on Promotion of Information and Communications Network Utilization and Information Protection), covering cybersecurity and data protection for online services [koreaNetworkAct]; and the Product Liability Act, which holds manufacturers accountable for damages arising from defects in AI-driven products, including software faults [koreaProductLiabilityAct]. Together, these regulations establish a robust, multi-layered foundation for trustworthy AI governance in South Korea.
@@ 4 Interpretation Challenges
    There are ”inherent concerns that the requirement for human oversight may be fundamentally incompatible with Agentic AI systems, which by definition are designed to act independently to achieve specific goals” [iapp2024agentic]. This creates a fundamental regulatory paradox where compliance requirements conflict with core system capabilities. The EU AI Act Article 14 [euaiact2024, Article 14] requires that ”high-risk AI systems shall be designed and developed in such a way that they can be effectively overseen by natural persons,” with oversight measures that are ”commensurate with the risks, level of autonomy and context of use”. However, the AI Act allows the Commission to consider ”the extent to which the AI system acts autonomously” when assessing risk, suggesting that more autonomous systems are more likely to be classified as ”high-risk” and thus require more oversight.
@@
    While the EU Code of Practice, Copyright Chapter, Measure 1.4 [eu2024codeofpractice] requires that Technical Model Documentation specify data collection sources, training datasets, and related information, it provides no corresponding infringement monitoring systems to verify compliance with these documentation requirements. The way that AI Office can monitor infringement cases is through:

### v2

@@ 3.8.3 South Korea
South Korea became the first jurisdiction outside the EU to enact a comprehensive AI law, the Basic Act on the Development of AI and the Establishment of Trust (“AI Basic Act”), officially passed on December 26, 2024, and slated to take effect in January 2026, following a one-year preparation period [34]. The Act focuses on “high-impact AI” systems deployed in critical areas such as healthcare, education, finance, employment, and essential public services. These systems must be designed to allow continuous, meaningful human oversight and intervention, and providers are required to notify users if they are interacting with AI-generated content or decisions, conduct risk assessments, maintain proper documentation, and implement effective control measures [34]. Additionally, foreign AI providers must appoint a local representative to handle regulatory communications [34].
This new law operates alongside existing statutes: the Personal Information Protection Act (PIPA), which mandates user consent and restricts the scope of personal data collection [30]; the Network Act (Act on Promotion of Information and Communications Network Utilization and Information Protection), covering cybersecurity and data protection for online services [29]; and the Product Liability Act, which holds manufacturers accountable for damages arising from defects in AI-driven products, including software faults [31]. Together, these regulations establish a robust, multi-layered foundation for trustworthy AI governance in South Korea.
@@ 4 Interpretation Challenges
    There are ”inherent concerns that the requirement for human oversight may be fundamentally incompatible with Agentic AI systems, which by definition are designed to act independently to achieve specific goals” [27]. This creates a fundamental regulatory paradox where compliance requirements conflict with core system capabilities. The EU AI Act Article 14 [18, Article 14] requires that ”high-risk AI systems shall be designed and developed in such a way that they can be effectively overseen by natural persons,” with oversight measures that are ”commensurate with the risks, level of autonomy and context of use”. However, the AI Act allows the Commission to consider ”the extent to which the AI system acts autonomously” when assessing risk, suggesting that more autonomous systems are more likely to be classified as ”high-risk” and thus require more oversight.
@@
    While the EU Code of Practice, Copyright Chapter, Measure 1.4 [10] requires that Technical Model Documentation specify data collection sources, training datasets, and related information, it provides no corresponding infringement monitoring systems to verify compliance with these documentation requirements. The way that AI Office can monitor infringement cases is through:
@@ 5 Conclusion
Acknowledgement
The authors would like to thank Jakub Łatkiewicz for his time and effort in reviewing this manuscript and for providing valuable comments and suggestions that helped improve its clarity and quality.